find two consecutive integers such that the sum of the first and four times the second is 119

Answer by checkley79(3341)

LET X & X+1 BE THE 2 INTEGERS.
X+4(X+1)=119
X+4X+4=119
5X=119-4
5X=115
X=115/5
X=23 ANS. FOR THE SMALLER INTEGER.
23+1=24 ANS. FOR THE GREATER INTEGER.
23+4*24=119
23+96=119
119=119